OpenStudy (anonymous):

simplify the radical expression. (square root of 10/121)

OpenStudy (campbell_st):

\[\sqrt{\frac{10}{121}} = \pm \frac{\sqrt{10}}{11}\]

OpenStudy (alexwee123):